OpenStudy (anonymous):

Solve the equation 5/4(a-8)= 2/3 a) 8 8/15 b) 8 2/3 c) 6 14/15 d) 7 1/15

OpenStudy (anonymous):

Hi I can help you

OpenStudy (anonymous):

Will you please?

OpenStudy (bohotness):

do you know how to work this out or slove it?

OpenStudy (anonymous):

not really. all i know is that reciprocal is used

OpenStudy (bohotness):

okay what do you think your answer is before we start?

OpenStudy (anonymous):

I honestly don't know. I was thinking probably b? 8 2/3?

OpenStudy (bohotness):

okay lets start Rearrange: Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 5/4*(a-8)-(2/3)=0

OpenStudy (bohotness):

got that part down?

OpenStudy (anonymous):

yes

OpenStudy (bohotness):

http://prntscr.com/6bfxyp

OpenStudy (anonymous):

okay

OpenStudy (bohotness):

Find the Least Common Multiple

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

but first wouldn't you do 5a-40/4 - 2/3? or would you do that after you find the LCM?

OpenStudy (bohotness):

:d

OpenStudy (bohotness):

he left denominator is : 4 The right denominator is : 3 Number of times each prime factor appears in the factorization of: Prime Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} 2 2 0 2 3 0 1 1 Product of all Prime Factors 4 3 12 Least Common Multiple: 12

OpenStudy (bohotness):

do you got that down?

OpenStudy (anonymous):

yes, i know how to do the LCM of fractions.

OpenStudy (bohotness):

Calculate multipliers for the two fractions <-- whats next

OpenStudy (bohotness):

Denote the Least Common Multiple by L.C.M Denote the Left Multiplier by Left_M Denote the Right Multiplier by Right_M Denote the Left Deniminator by L_Deno Denote the Right Multiplier by R_Deno Left_M = L.C.M / L_Deno = 3 Right_M = L.C.M / R_Deno = 4

OpenStudy (anonymous):

ok

OpenStudy (bohotness):

http://prntscr.com/6bg84r